# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

### v1

@@ Abstract
Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University,4Imperial College London 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
@@
1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (Waller and Anderson, 2021; Iandoli et al., 2021). It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (Garimella, 2018). Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence. As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
@@ 1 Introduction
    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping, deindividuation, extreme language)?
We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context. Extensive qualitative analysis reveals how event salience, linguistic norms, and platform affordances shape polarization dynamics across languages. The complete pipeline can be found in Figure 1. We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
@@ 2 Related Work
Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (Waller and Anderson, 2021; Iandoli et al., 2021; Garimella, 2018). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance. Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (Arora et al., 2022), and has highlighted its relationship with hate speech, fragmentation, and incivility (Bozdag and van den Hoven, 2020; Mathew et al., 2021).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (Kubin and von Sikorski, 2021; Barberá, 2020; Gitlin, 2016; Soares and Recuero, 2021). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (Kubin and von Sikorski, 2021). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (Basile and others, 2019; Mohammad and others, 2021; Pamungkas et al., 2020; Ousidhoum et al., 2024) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
@@ 3.1 Data Collection
We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums. We use a dynamic keyword-driven strategy tailored for each language. Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events. Table 1 shows the languages covered, data splits and total number of instances annotated for each language.
@@
Language Source(s) Train Dev Test Total
Amharic Facebook, X 3500 500 1000 5000
Arabic Facebook, X, Threads, News 1482 212 424 2118
English X, BlueSky, News 2117 303 605 3025
German X, BlueSky, Reddit 2426 347 694 3467
Hausa Facebook , X 3893 557 1113 5563
Spanish X, BlueSky 1400 201 401 2002
Urdu X 1960 280 560 2800
Table 1: Dataset sources and split sizes for all three tasks as per each language.
@@ 3.2 Annotation Process
Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators.
Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. Annotators were instructed to:
@@
  • If the text is classified as polarized, tag its manifestations (stereotyping, vilification, dehumanization, deindividuation, extreme language, lack of empathy, invalidation).
@@
Annotator Recruitment and Annotation Process
We used both expert and crowd-sourced annotation strategies. Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific. To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa. Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
@@
  • Amharic: We used train annotators to annotated 5,000 samples in five batches using the POTATO annotation tool. Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control. The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
Figure 3: Percentage of polarization per language
(a) Types
(b) Manifestation
Figure 4: Annotation statistics for polarization classification, types, and manifestations
3.3 Dataset Characteristics
Polarization Prevalence: As shown in Figure 3, Amharic (75%) and Urdu (66%) has the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate. Spanish (53%), German (46%), and English (34%) fall in the mid-range. Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects. Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression. Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization. Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
Types of Polarization: Figure 4 show types of polarization. Political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%). Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German. Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%). Thematic emphasis (e.g., migration, elections) shapes type distribution. Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
Polarization Manifestations: Figure 4 show types of Manifestation. Stereotyping and vilification are most frequent in Urdu, Amharic, German, and Spanish. Urdu’s vilification rate is highest (55%). Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%). Invalidation and dehumanization peak in German and Urdu. Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations. Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization. The POLAR dataset (see Table 2 for detailed statistics of POLAR dataset) offers a robust benchmark for examining these effects at scale.
@@ 4.1 Experimental Setup
  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., incivility, stereotyping, dehumanization).
For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table 1. We pursued two main experimental paradigms:
@@
    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023) for monolingual and crosslingual experiments.
@@
    Evaluating Large Language Models (LLMs) in Zero- and Few-shot Settings: We tested models including QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, and GPT4o/mini.
@@ 4.2.1 Monolingual Setup Results
To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023). Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
@@ 4.2.3 Zero and Few-shot LLM Performance
We evaluated five state-of-the-art LLMs (QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, GPT4o/mini) in a zero- and few-shot setting across languages and for the first ask. Larger models such as GPT4o and GPT4o-mini outperformed smaller ones, especially in capturing subtle polarization cues.
GPT 4o GPT 4o-mini Mistral-7B Llama-3.1-8B Qwen3-8B
Zero-shot Few-shot Zero-shot Few-shot Zero-shot Zero-shot Zero-shot
Amharic 71.30 68.74 55.40 62.59 71.92 36.32 58.09
Arabic 71.51 79.65 58.21 75.95 33.65 34.82 54.54
English 75.46 79.56 64.78 80.93 50.34 56.75 72.07
German 72.04 71.12 63.03 72.44 52.63 56.22 61.67
Hausa 34.73 44.87 28.21 44.26 23.42 18.51 25.06
Spanish 72.69 65.76 62.51 69.00 56.31 41.75 61.11
Urdu 72.25 78.83 59.19 74.66 70.47 67.26 69.63

### v2

@@ Abstract
Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University, 4Imperial College London, 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (waller2021quantifying; iandoli2021impact). It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (garimella2018polarization). Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence. As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
@@ 1 Introduction
    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping/deindividuation, vilification, dehumanization, extreme language)?
We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context. The complete pipeline can be found in Figure 1. We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
@@ 2 Related Work
Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (waller2021quantifying; iandoli2021impact; garimella2018polarization). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance. Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (arora2022polarization), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew2021HateXplain).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (kubin2021role; barbera2020social; gitlin2016outrage; soares2021hashtag). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (kubin2021role). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (basile2019semeval; pamungkas2020misogynistic; ousidhoum2024semeval) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
@@ 3 POLAR Dataset Construction
We aim to construct a multilingual, multi-event, and multi-platform social media dataset spanning multiple continents. The dataset is designed to cover Indo-European and Afro-Asiatic language families, with a particular focus on balancing linguistic diversity and resource availability. It includes high-resource languages such as English, Spanish, and German; medium-resource languages like Arabic and Urdu; and low-resource languages such as Hausa and Amharic. While several Indo-European languages are represented (e.g., English, Spanish, German, Urdu), the dataset also incorporates key Afro-Asiatic languages (e.g., Hausa, Arabic, Amharic) to ensure broader geographical and linguistic representation. The following section will discuss the dataset construction in detail.
3.1 Operational Definitions
In this study, polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups. We also classify the types of polarization as follows:
  • Political/ideological polarization: This type of extremism focuses on division, intolerance, and conflict between political parties and followers.
  • Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
  • Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers
  • Gender polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender.
  • Sexual orientation polarization: This refers to the increasing division and distinction between different sexual orientations within society, often leading to heightened tensions, misunderstandings, conflicts, or marginalization among various groups.
  • Other: Polarized texts targeting other groups/identities such as economy class, technology, media, polarization, etc are classified as others.
Furthermore, we define a polarized text manifestations or tactics if it exhibits one or more of the following characteristics:
  • Stereotype: A generalized belief that attributes specific characteristics to all members of a group, often neglecting individual differences, thereby reducing complex personalities to simplistic and uniform representations.
  • Vilification: The act of defaming or demonizing a particular group, person, or entity by inciting fear, often through exaggeration, misrepresentation, or biased framing that portrays the subject negatively and harmfully.
  • Dehumanization: The process of depriving a group or individual of their human qualities or personality by comparing them to animals, machines, or objects, or otherwise denying their humanity, dignity, or individuality.
  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like "always," "never," "worst," or "best," and presenting issues in a dichotomous manner such as "us versus them" or "right versus wrong.
  • Lack of Empathy or Understanding: The absence of compassion or recognition for other viewpoints or experiences in the text.
  • Invalidation: The act of denying or dismissing the identity and existence of individuals or groups, thereby rejecting their sense of self and their presence. See Appendix A for further details and examples for each manifestation of polarization.
@@ 3.2 Data Collection
We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums. We use a dynamic keyword-driven strategy tailored for each language. Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events. Table˜1 illustrates the compositions, size, and event coverage of the dataset.
@@
Language Source(s) Event Train Dev Test Total Instances
Amharic Facebook, X Tigray War in Northern Ethiopia 3500 500 1000 5000
Arabic Facebook, X, Threads, News Social and economic issues 1482 212 424 2118
English X, BlueSky, News US elections and international conflicts 2117 303 605 3025
German X, BlueSky, Reddit Election discourse and migration debates 2426 347 694 3467
Hausa Facebook, X Religious and ideological discussions in Africa 3893 557 1113 5563
Spanish X, BlueSky Abortion, migration, and indigenous/gender rights 1400 201 401 2002
Urdu X Political and sectarian conflicts 1960 280 560 2800
Table 1: Summary of dataset composition: sources, sizes, and event coverage by task and language.
@@ 3.3 Annotation Process
Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators. For crowdsourced annotators, we first filtered candidates based on their previous annotation experience and annotation agreement. Specifically, we selected those who had completed prior tasks with an agreement score of at least 0.8. We also conducted a pilot task consisting of 50 items to further filter annotators based on their agreement scores and average labeling time. For community annotators, we provided a training session followed by a pilot task to assess their performance before assigning the next batch of annotation tasks. During the annotation process, we continuously monitored annotator performance in a batch-wise manner to ensure that the overall process remained on track.
Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. The English annotation guidelines are provided in Appendix A. Annotators were instructed to:
@@
  • If the text is classified as polarized, tag its manifestations (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
@@
Annotator Recruitment and Annotation Process: We used both expert and crowd-sourced annotation strategies. Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific. To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa. Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
  • Amharic: We used trained annotators to annotated 5,000 samples in five batches using the POTATO annotation tool. Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control. The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
Figure 3: Percentage of polarization per language
3.4 Dataset Characteristics
@@
We now analyze key trends and insights across the dataset based on annotated labels. Table 2 provides a detailed quantitative breakdown of positive labels for each annotation task.
Polarization Prevalence: As shown in Figure 3 and further detailed in Table 2, Amharic (75%) and Urdu (66%) have the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate. Spanish (53%), German (46%), and English (34%) fall in the mid-range. Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects. Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression. Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization. Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
Types of Polarization: As illustrated in Figure LABEL:fig:types and Table 2, political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%). Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German. Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%). Thematic emphasis (e.g., migration, elections) shapes type distribution. Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
Polarization Manifestations: Figure LABEL:fig:manifestation and Table 2 show the types of manifestations. Stereotyping/deindividuation and vilification are most frequent in Urdu, Amharic, German, and Spanish. Urdu’s vilification rate is highest (55%). Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%). Invalidation and dehumanization peak in German and Urdu. Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations. Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization. The POLAR dataset offers a robust benchmark for examining these effects at scale.
@@ 4.1 Experimental Setup
  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., stereotyping, vilification, dehumanization).
For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜1. We pursued two main experimental paradigms:
@@
    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3) for monolingual and crosslingual experiments.
@@
[…]
@@ 4.2.1 Monolingual Setup Results
To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3). Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
@@ 4.2.3 Zero and Few-shot LLM Performance
We evaluated seven state-of-the-art LLMs (Qwen3-8B, LLaMA-3.1-8B, Mixtral-8x7B, GPT-4o, GPT-4o-mini, GPT-4.1-mini, and Llama-4-Scout-17B-16E-Instruct) across languages. The prompts used for evaluation are provided in the Appendix B. We explicitly provided definitions of tasks and leveraged the models’ chain-of-thoughts capabilities by encouraging reasoning before labeling. In the few-shot setting, we provided 4 example instances for each task. To assess the impact of prompt language, we also compared model performance when prompted in English versus the respective target language.
As Table  4 illustrates, the closed-source models GPT-4o, GPT-4o mini, and GPT-4.1 mini, consistently outperform the other four LLMs, demonstrating superior performance in polarization recognition. However, their performance on Hausa remains notably low, highlighting both the challenge of understanding low-resource languages and the value of our multilingual dataset in exposing the limitations.
In contrast, smaller open-source models, such as Qwen3-8B, LLaMA-3.1-8B, and Mixtral-8x7B, achieve lower F1-Macro scores and perform inconsistently across languages. Their F1-Macro scores improve by 5 10% when provided with few-shot examples. However, the larger LLaMA model, LLaMA-4-scout-17B, demonstrates performance comparable to GPT-4o in polarization classification.
An interesting phenomenon is observed: Smaller models show improved classification performance in the few-shot setting, while larger models such as the GPT series and LLaMA-4-17B do not consistently gain additional improvement with few-shot examples. This suggests that few-shot learning offers advantages to smaller models than to larger ones.
Task Lang GPT 4o GPT 4o-mini GPT 4.1-mini Mistral-7B Qwen3-8B Llama-3.1-8B Llama4-17B
Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot
1. Polarization Amharic 71.30 68.74 55.40 58.04 63.52 63.20 44.84 48.24 55.60 65.51 50.98 58.84 69.96 67.91
Arabic 71.51 79.65 58.21 68.41 76.20 79.58 33.65 53.06 65.95 70.76 61.51 67.21 73.58 77.45
English 75.46 79.56 64.78 80.30 77.43 77.48 53.78 70.00 69.09 66.86 66.55 69.90 76.90 75.48
German 72.04 71.12 63.03 68.49 73.77 69.60 56.17 61.77 65.42 68.76 59.90 65.40 70.89 70.35
Hausa 34.73 44.87 28.21 34.07 39.45 48.06 19.82 28.15 45.91 49.18 38.73 40.80 47.85 54.53
Spanish 72.69 65.76 62.51 67.96 71.28 69.22 51.24 61.00 65.42 68.76 60.20 62.53 66.11 66.27
Urdu 72.25 78.83 59.19 69.92 73.15 72.63 48.27 54.79 66.99 70.89 55.61 59.77 73.93 68.50
2. Types Amharic 39.67 36.59 33.8 34.89 32.49 34.44 44.21 63.24 63.07 75.90 78.26
Arabic 35.00 39.30 37.18 36.82 33.07 36.77 49.39 53.91 53.22 52.86 55.03
English 34.96 35.67 33.79 38.73 35.22 38.22 48.28 68.50 67.37 62.71 67.09
German 58.86 58.58 54.36 57.94 52.70 55.15 44.10 71.82 70.62 73.89 76.20
Hausa 18.96 19.11 16.90 18.15 17.00 18.86 43.34 67.24 70.47 71.75 74.66
Spanish 48.98 52.93 52.82 55.09 48.89 49.76 43.15 65.96 67.87 69.82 71.43
Urdu 55.6 59.64 53.97 54.78 51.82 55.09 48.40 63.36 64.83 60.01 61.52
3. Manifestations Amharic 44.09 45.59 40.01 44.78 40.53 42.49 0 0 0 0 0
Arabic 51.90 56.90 46.96 51.01 52.57 60.76 0 0 0 0 0
English 22.95 25.87 20.43 25.78 28.41 26.67 0 0 0 0 0
German 42.84 45.16 43.33 53.78 37.62 41.23 0 0 0 0 0
Hausa 8.58 8.14 8.42 8.37 8.55 9.33 0 0 0 0 0
Spanish 41.00 40.99 39.85 46.33 38.82 40.66 0 0 0 0 0
Urdu 52.89 54.43 48.25 58.17 48.34 51.89 0 0 0 0 0